# BayTTA: Uncertainty-aware medical image classification with optimized test-time augmentation using Bayesian model averaging

### v1

@@ Abstract
Test-time augmentation (TTA) is a well-known technique employed during the testing phase of computer vision tasks. It involves aggregating multiple augmented versions of input data. Combining predictions using a simple average formulation is a common and straightforward approach after performing TTA. This paper introduces a novel framework for optimizing TTA, called BayTTA (Bayesian-based TTA), which is based on Bayesian Model Averaging (BMA). First, we generate a model list associated with different variations of the input data created through TTA. Then, we use BMA to combine model predictions weighted by their respective posterior probabilities. Such an approach allows one to take into account model uncertainty, and thus to enhance the predictive performance of the related machine learning or deep learning model. We evaluate the performance of BayTTA on various public data, including three medical image datasets comprising skin cancer, breast cancer, and chest X-ray images and two well-known gene editing datasets, CRISPOR and GUIDE-seq. Our experimental results indicate that BayTTA can be effectively integrated into state-of-the-art deep learning models used in medical image analysis as well as into some popular pre-trained CNN models such as VGG-16, MobileNetV2, DenseNet201, ResNet152V2, and InceptionRes-NetV2, leading to the enhancement in their accuracy and robustness performance. The source code of the proposed BayTTA method is freely available at: https://github.com/Z-Sherkat/BayTTA.
keywords:
MSC:
[…]
journal: Medical Image Analysis
@@
Fig. 1: Examples of original benign and malignant skin cancer, breast cancer, and chest X-ray images and their augmented versions considered in our study.
1 Introduction
@@
Deep learning models are highly effective tools for identifying complex patterns in data. However, they are also prone to overfitting and memorizing the training data rather than generalizing real data patterns. Data augmentation is a critical strategy to overcome this issue, which involves synthesizing new training samples by applying to the original data a series of transformations and perturbations, such as random erasing, flipping, rotating, and scaling (see Fig. 1). Leveraging data augmentation techniques effectively introduces variability in the training data, leading to a more robust and general model. Machine learning (ML) and deep learning (DL) models equipped with data augmentation can handle the complexities inherent in medical image data, often leading to an enhanced diagnostic accuracy ([56, 8, 21, 9]).
@@
Test-Time Augmentation (TTA) is a popular data augmentation technique which can increase the prediction power of ML and DL models. TTA involves applying data augmentation on test input rather than on training data. It provides a more reliable estimate of the target variable by averaging the predictions across augmented inputs to find the output result. While data augmentation is mainly used for improving training data diversity to enhance the model generalization on unseen data, TTA produces multiple versions of each test image to obtain a more robust estimate of the target variable ([32, 40]).
@@
Recently, Uncertainty Quantification (UQ) methods have been effectively applied in a variety of research domains, contributing to the robustness and reliability of the optimization and decision-making processes. There are two primary categories of uncertainty ([1, 30]), referred to as aleatoric and epistemic. While aleatoric uncertainty pertains to the inherent characteristic of the data distribution, epistemic uncertainty relates to the lack of knowledge about the optimal model. Uncertainty quantification is a vital tool for addressing the limitations of data (aleatoric uncertainty) and models (epistemic uncertainties) in various scientific and engineering applications, and thus for improving the trustworthiness of predictions, especially in the testing phase ([3]). Although deep learning models often outperform traditional machine learning methods, overconfident predictions remain their crucial issue, leading to excessive prediction errors ([51]). By combining uncertainty quantification with TTA, we can gain a deeper insight into the reliability and robustness of model predictions ([57, 6, 5]). This is particularly beneficial for machine learning and deep learning applications in healthcare, autonomous systems, and safety-critical domains, where understanding and managing uncertainty may be critical.
@@
  • 3.
    We assess the performance of BayTTA on three public medical image datasets, including skin cancer, breast cancer, and chest X-ray images available on the Kaggle and Mendeley data repositories as well as on two popular gene editing datasets, CRISPOR and GUIDE-seq, generated using the CRISPR-Cas9 technology ([23]). We demonstrate that our method can be applied successfully with different deep learning network architectures. Our results indicates a superior performance of the proposed method compared to standard averaging.
  • 4.
@@
2 Related work
2.1 Uncertainty quantification
@@
Technical progress: Uncertainty quantification (UQ) is essential to enhance the credibility of predictions during the testing phase. Since the excessive confidence of deep neural networks may lead to prediction errors, it is imperative to address the issue of overconfident predictions in order to improve their reliability and trustworthiness ([24]). As deep learning models are now constantly used in critical areas, the ability to quantify and manage uncertainty becomes increasingly vital ([29, 41, 2]).
@@
Nowadays, UQ has important applications in image processing, computer vision, medical image analysis, diagnostic modeling, and healthcare decision-making ([1, 37, 38, 48]). There are several well-known methods for measuring uncertainty such as Monte Carlo dropout ([20]), Variational Inference ([59, 47]), Deep Ensembles ([13, 46, 4]), and Bayesian Deep Ensembles ([26]). Bayesian Model Averaging (BMA) is another effective technique used to take into account prediction uncertainty ([60, 42, 31, 7]). Development of BMA in the context of model uncertainty has been influenced by the seminal works of [15, 16] and [28] as their original methods provide insight into the quantification of both epistemic and aleatoric uncertainties.
@@
UQ application in medical image analysis: Nowadays, UQ methods have become a tool of choice for estimating the uncertainty associated with disease detection, diagnosis, medical image segmentation, and identification of the region of interest (ROI) in medical image analysis. In their recent work, [3] have introduced an efficient hybrid dynamic model of uncertainty quantification, called TWDBDL. The model is based on the Three-Way Decision (TWD) theory and Bayesian Deep Learning (BDL) methods used together to improve the trustworthiness of predictions in skin cancer detection. [17] used a probabilistic variational autoencoders (VAEs), i.e. a Monte Carlo technique to generate pixel uncertainty maps, and Stein’s Unbiased Risk Estimator (URE) to provide accurate uncertainty estimations in knee magnetic resonance imaging. [22] introduced the UA-ConvNet, i.e. an uncertainty-aware Convolutional Neural Network (CNN) model, for COVID-19 detection in chest X-ray (CXR) images. The model estimates the uncertainty based on the EfficientNet-B3 Bayesian network supplemented with Monte Carlo dropout.
@@
[41] designed a novel web server, called DUNEScan, for uncertainty estimation in CNN models applied to skin cancer detection. The web server employs binary dropout to compare the average model predictions, providing visualization for uncertainty to diagnose precisely skin cancer cases. [25] proposed a novel dynamic multi-scale convolutional neural network (DM-CNN) that leverages a hierarchical dynamic uncertainty quantification attention (HDUQ-Attention) submodel. HDUQ-Attention includes a tuning block for adjusting the attention weights as well as Monte Carlo dropout for quantifying uncertainty. The experiments conducted on skin disease images (HAM10000), colorectal cancer images (NCT-CRC-HE-100K), and lung disease images (OCT2017 and Chest X-ray) demonstrated that the DM-CNN model accurately quantifies uncertainty, showing a stable performance.
@@
UQ application in gene editing: Uncertainty quantification can be effectively applied in the context of gene editing in order to improve the trustworthiness of on- and off-target predictions in CRISPR-Cas9 experiments ([50]).
@@
For example, [62] presented a deep learning model for off-target activity prediction, employing data augmentation to mitigate the class imbalance issue. The authors collected data from two source types, i.e. in vitro and cell-based experiments, to increase the size of the positive class samples (off-targets). They suggested synthetically expanding the number of positive samples by rotating the sgRNA-DNA encoded images by 90, 180, and 270 degrees, respectively, to enhance the model competency.
@@
Moreover, [35] have recently introduced a pioneering method that incorporates uncertainty into off-target predictions. Their approach offers an interpretable evaluation of Cas9-gRNA and Cas12a-gRNA specificity through deep kernel learning, estimating a gRNA’s cleavage efficiency with a corresponding confidence interval.
Fig. 2: A schematic view of a conventional test-time augmentation (TTA) process.
2.2 Test-time augmentation
@@
Technical progress: Test-time augmentation (TTA) is a well-known technique that applies data augmentation during the testing phase. TTA has multiple benefits, including improving the model generalization and reliability capacity ([53, 12]), estimating uncertainty in model predictions ([57, 6, 5]), and boosting accuracy in classification and segmentation tasks ([36, 54, 27, 49]). Figure 2 presents a detailed flowchart of a typical TTA process.
@@
Recently, some advanced TTA methods using diverse aggregation strategies have been proposed. They include, among others, selective augmentation techniques such as the instance-aware TTA based on a loss predictor ([34]), the instance-level TTA with entropy weight method ([11]), and the selective-TTA method ([52]).
@@
[39] presented a straightforward and efficient method based on a policy search algorithm, known as greedy policy search (GPS), designed to optimize image classification. The method aims to optimize and determine the most effective data augmentation strategy to be applied during the testing phase of a machine learning process, and thus to improve the prediction accuracy and robustness of ML models. [34] developed an instance-aware test-time augmentation approach that employs a loss predictor to dynamically select test-time transformations based on the expected losses for individual instances. Furthermore, [11] introduced an instance-level TTA with Entropy Weight Method (EWM) as an innovative approach to improve the accuracy and robustness of classification models.
@@
The paper of [49] describes a new method for aggregating model predictions obtained from TTA. Unlike the traditional approach of averaging model predictions, this method focuses on learning different augmentation weights to aggregate predictions obtained from transformations during TTA. The authors noticed that existing aggregation methods, based on the mean or the maximum of predictions obtained from augmented images, may not be optimal because they do not consider the relationship between the original test image and its augmented versions. The paper also offers insights into the point when TTA can be helpful and provides guidance regarding the use of different TTA policies. Moreover, Shanmugam and co-authors characterized the cases where TTA transforms correct predictions into incorrect ones, and vice versa.
@@
TTA application in medical image analysis: TTA has been extensively studied by the medical imaging community due to its capability to contribute to model robustness and improve the trustworthiness and generalization of predictions during the testing phase. [57] conducted critical research in the context of deep convolutional neural network (CNN)-based medical image segmentation. Their work focuses on epistemic and aleatoric uncertainty analysis at pixel and structure levels, providing valuable insights into the reliability of segmentation results. [19] introduced a simple and powerful task-agnostic semi-supervised active learning segmentation approach, called TAAL (TTA for Active Learning). TAAL exploits unlabeled samples during training and sampling phases by using a technique known as cross-augmentation consistency.
3 Methodology
@@
In this section, we elaborate on the background, foundation, and comprehensive explanation of the methodological procedures used within the proposed method. Our new method, called BayTTA, aims to optimize TTA using BMA and uncertainty estimation. The first component of our method involves formulating a mathematical model that applies BMA to output predictions generated from multiple transformed versions of the input data (e.g. input images). The second component of the method estimates the uncertainty of the predictions obtained from an augmented set of test images associated with image transformations or noise. Together, these components offer a comprehensive approach to enhance the robustness and reliability of ML or DL models during the testing phase. An overview of the proposed BayTTA method is presented in Figure 3.
@@
Fig. 3: An overview of the proposed BayTTA method. During the testing phase: (a) TTA generates a model list from a set of fixed augmented images, and (b) BMA is then applied to combine and aggregate the predictions from the model list.
@@
Traditional TTA involves applying a number of data augmentation techniques, including rotation, cropping, flipping, and brightness adjustments, to the test data before making predictions. This widely-used approach leverages ensemble methods, such as averaging or taking the majority vote of predictions made on augmented samples, to reduce the impact of random noise and variation in the test data. TTA usually produces stable and accurate predictions by averaging multiple augmented versions of the test dataset ([57, 58]).
@@
Our study also implements BMA to integrate predictions from the model list, thus deviating from the conventional averaging approach used in TTA, to yield more precise and robust outcomes. Let us focus on a single image $I_{y}$ from the test dataset, where $y\in\{0,1\}$ denotes its classification label. In a scenario generating $k$ augmented versions of the test image in TTA, the input to the BMA comes from the vector ${\mathbf{x}=\{x_{1},x_{2},...,x_{k+1}\}}$, where each element corresponds to a prediction from the TTA procedure. Here, $x_{1}$ denotes the predictor variable derived from the original test image $I_{y}$ and $\{x_{2},x_{3},...,x_{k+1}\}$ denote the predictor variables from augmented versions of this image.
@@
BMA involves generating various models by combining predictor variables and selecting candidate models based on likelihood. To formalize this process, we introduce $\mathcal{P(\mathbf{x})}$, a non-empty power set of $\mathbf{x}$, representing all combinations of predictor variables. Subsequently, utilizing $I\ \in\mathcal{P(\mathbf{x})}$, we select desired predictor variables from $\mathbf{x}$ for model $M_{I}$, where $\mathbf{x}_{I}$ represents the input of the respective model $M_{I}$. In the context of Bayesian model averaging logistic regression, we further elucidate the mathematical formulations that define our algorithm in the subsequent discussion.
@@
4 Experimental evaluations and discussion

### v2

@@ Abstract
Test-time augmentation (TTA) is a well-known technique employed during the testing phase of computer vision tasks. It involves aggregating multiple augmented versions of input data. Combining predictions using a simple average formulation is a common and straightforward approach after performing TTA. This paper introduces a novel framework for optimizing TTA, called BayTTA (Bayesian-based TTA), which is based on Bayesian Model Averaging (BMA). First, we generate a prediction list associated with different variations of the input data created through TTA. Then, we use BMA to combine predictions weighted by the respective posterior probabilities. Such an approach allows one to take into account model uncertainty, and thus to enhance the predictive performance of the related machine learning or deep learning model. We evaluate the performance of BayTTA on various public data, including three medical image datasets comprising skin cancer, breast cancer, and chest X-ray images and two well-known gene editing datasets, CRISPOR and GUIDE-seq. Our experimental results indicate that BayTTA can be effectively integrated into state-of-the-art deep learning models used in medical image analysis as well as into some popular pre-trained CNN models such as VGG-16, MobileNetV2, DenseNet201, ResNet152V2, and InceptionRes-NetV2, leading to the enhancement in their accuracy and robustness performance. The source code of the proposed BayTTA method is freely available at: https://github.com/Z-Sherkat/BayTTA.
Index Terms:
Bayesian Model Averaging, Deep Learning, Test-time Augmentation, Uncertainty Quantification
@@
Figure 1: Examples of original benign and malignant skin cancer, breast cancer, and chest X-ray images and their augmented versions considered in our study.
I Introduction
@@
Deep learning models are highly effective tools for identifying complex patterns in data. However, they are also prone to overfitting and memorizing the training data rather than generalizing real data patterns. Data augmentation is a critical strategy to overcome this issue, which involves synthesizing new training samples by applying to the original data a series of transformations and perturbations, such as random erasing, flipping, rotating, and scaling (see Fig. 1). Leveraging data augmentation techniques effectively introduces variability in the training data, leading to a more robust and general model. Machine learning (ML) and deep learning (DL) models equipped with data augmentation can handle the complexities inherent in medical image data, often leading to an enhanced diagnostic accuracy [1, 2, 3, 4].
@@
Test-Time Augmentation (TTA) is a popular data augmentation technique which can increase the prediction power of ML and DL models. TTA involves applying data augmentation on test input rather than on training data. It provides a more reliable estimate of the target variable by averaging the predictions across augmented inputs to find the output result. While data augmentation is mainly used for improving training data diversity to enhance the model generalization on unseen data, TTA produces multiple versions of each test image to obtain a more robust estimate of the target variable [5, 6].
@@
Recently, Uncertainty Quantification (UQ) methods have been effectively applied in a variety of research domains, contributing to the robustness and reliability of the optimization and decision-making processes. There are two primary categories of uncertainty [7, 8], referred to as aleatoric and epistemic. While aleatoric uncertainty pertains to the inherent characteristic of the data distribution, epistemic uncertainty relates to the lack of knowledge about the optimal model. Uncertainty quantification is a vital tool for addressing the limitations of data (aleatoric uncertainty) and models (epistemic uncertainties) in various scientific and engineering applications, and thus for improving the trustworthiness of predictions, especially in the testing phase [9]. Although deep learning models often outperform traditional machine learning methods, overconfident predictions remain their crucial issue, leading to excessive prediction errors [10]. By combining uncertainty quantification with TTA, we can gain a deeper insight into the reliability and robustness of model predictions [11, 12, 13]. This is particularly beneficial for machine learning and deep learning applications in healthcare, autonomous systems, and safety-critical domains, where understanding and managing uncertainty may be critical.
@@
  • We assess the performance of BayTTA on three public medical image datasets, including skin cancer, breast cancer, and chest X-ray images available on the Kaggle and Mendeley data repositories as well as on two popular gene editing datasets, CRISPOR and GUIDE-seq, generated using the CRISPR-Cas9 technology [14]. We demonstrate that our method can be applied successfully with different deep learning network architectures. Our results indicates a superior performance of the proposed method compared to standard averaging.
@@
II Related work
II-A Uncertainty quantification
@@
Technical progress: Uncertainty quantification (UQ) is essential to enhance the credibility of predictions during the testing phase. Since the excessive confidence of deep neural networks may lead to prediction errors, it is imperative to address the issue of overconfident predictions in order to improve their reliability and trustworthiness [15]. As deep learning models are now constantly used in critical areas, the ability to quantify and manage uncertainty becomes increasingly vital [16, 17, 18].
@@
Nowadays, UQ has important applications in image processing, computer vision, medical image analysis, diagnostic modeling, and healthcare decision-making [7, 19, 20, 21]. There are several well-known methods for measuring uncertainty such as Monte Carlo dropout [22], Variational Inference [23, 24], Deep Ensembles [25, 26, 27], and Bayesian Deep Ensembles [28]. Bayesian Model Averaging (BMA) is another effective technique used to take into account prediction uncertainty [29, 30, 31, 32]. Development of BMA in the context of model uncertainty has been influenced by the seminal works of [33, 34] and [35] as their original methods provide insight into the quantification of both epistemic and aleatoric uncertainties.
@@
UQ application in medical image analysis: Nowadays, UQ methods have become a tool of choice for estimating the uncertainty associated with disease detection, diagnosis, medical image segmentation, and identification of the region of interest (ROI) in medical image analysis. In their recent work, [9] have introduced an efficient hybrid dynamic model of uncertainty quantification, called TWDBDL. The model is based on the Three-Way Decision (TWD) theory and Bayesian Deep Learning (BDL) methods used together to improve the trustworthiness of predictions in skin cancer detection. Edupuganti et al. [36] used a probabilistic variational autoencoders (VAEs), i.e. a Monte Carlo technique to generate pixel uncertainty maps, and Stein’s Unbiased Risk Estimator (URE) to provide accurate uncertainty estimations in knee magnetic resonance imaging. Gour et al. [37] introduced the UA-ConvNet, i.e. an uncertainty-aware Convolutional Neural Network (CNN) model, for COVID-19 detection in chest X-ray (CXR) images. The model estimates the uncertainty based on the EfficientNet-B3 Bayesian network supplemented with Monte Carlo dropout.
@@
Mazoure et al. [17] designed a novel web server, called DUNEScan, for uncertainty estimation in CNN models applied to skin cancer detection. The web server employs binary dropout to compare the average model predictions, providing visualization for uncertainty to diagnose precisely skin cancer cases. Han et al. [38] proposed a novel dynamic multi-scale convolutional neural network (DM-CNN) that leverages a hierarchical dynamic uncertainty quantification attention (HDUQ-Attention) submodel. HDUQ-Attention includes a tuning block for adjusting the attention weights as well as Monte Carlo dropout for quantifying uncertainty. The experiments conducted on skin disease images (HAM10000), colorectal cancer images (NCT-CRC-HE-100K), and lung disease images (OCT2017 and Chest X-ray) demonstrated that the DM-CNN model accurately quantifies uncertainty, showing a stable performance.
@@
UQ application in gene editing: Uncertainty quantification can be effectively applied in the context of gene editing in order to improve the trustworthiness of on- and off-target predictions in CRISPR-Cas9 experiments [39].
@@
For example, Zhang et al. [40] presented a deep learning model for off-target activity prediction, employing data augmentation to mitigate the class imbalance issue. The authors collected data from two source types, i.e. in vitro and cell-based experiments, to increase the size of the positive class samples (off-targets). They suggested synthetically expanding the number of positive samples by rotating the sgRNA-DNA encoded images by 90, 180, and 270 degrees, respectively, to enhance the model competency.
@@
Moreover, Kirillov et al. [41] have recently introduced a pioneering method that incorporates uncertainty into off-target predictions. Their approach offers an interpretable evaluation of Cas9-gRNA and Cas12a-gRNA specificity through deep kernel learning, estimating a gRNA’s cleavage efficiency with a corresponding confidence interval.
II-B Test-time augmentation
@@
Technical progress: Test-time augmentation (TTA) is a well-known technique that applies data augmentation during the testing phase. TTA has multiple benefits, including improving the model generalization and reliability capacity [42, 43], estimating uncertainty in model predictions [11, 12, 13], and boosting accuracy in classification and segmentation tasks [44, 45, 46, 47]. Fig. 2 presents a detailed flowchart of a typical TTA process.
Figure 2: A schematic view of a conventional test-time augmentation (TTA) process.
@@
Recently, some advanced TTA methods using diverse aggregation strategies have been proposed. They include, among others, selective augmentation techniques such as the instance-aware TTA based on a loss predictor [48], the instance-level TTA with entropy weight method [49], and the selective-TTA method [50].
@@
[51] presented a straightforward and efficient method based on a policy search algorithm, known as greedy policy search (GPS), designed to optimize image classification. The method aims to optimize and determine the most effective data augmentation strategy to be applied during the testing phase of a machine learning process, and thus to improve the prediction accuracy and robustness of ML models. Kim et al. [48] developed an instance-aware test-time augmentation approach that employs a loss predictor to dynamically select test-time transformations based on the expected losses for individual instances. Furthermore, Chun et al. [49] introduced an instance-level TTA with Entropy Weight Method (EWM) as an innovative approach to improve the accuracy and robustness of classification models.
@@
The paper of D. Shanmugam [47] describes a new method for aggregating model predictions obtained from TTA. Unlike the traditional approach of averaging model predictions, this method focuses on learning different augmentation weights to aggregate predictions obtained from transformations during TTA. The authors noticed that existing aggregation methods, based on the mean or the maximum of predictions obtained from augmented images, may not be optimal because they do not consider the relationship between the original test image and its augmented versions. The paper also offers insights into the point when TTA can be helpful and provides guidance regarding the use of different TTA policies. Moreover, Shanmugam and co-authors characterized the cases where TTA transforms correct predictions into incorrect ones, and vice versa.
@@
TTA application in medical image analysis: TTA has been extensively studied by the medical imaging community due to its capability to contribute to model robustness and improve the trustworthiness and generalization of predictions during the testing phase. Wang et al. [11] conducted critical research in the context of deep convolutional neural network (CNN)-based medical image segmentation. Their work focuses on epistemic and aleatoric uncertainty analysis at pixel and structure levels, providing valuable insights into the reliability of segmentation results. Gaillochet et al. [52] introduced a simple and powerful task-agnostic semi-supervised active learning segmentation approach, called TAAL (TTA for Active Learning). TAAL exploits unlabeled samples during training and sampling phases by using a technique known as cross-augmentation consistency.
III Methodology
@@
In this section, we elaborate on the background, foundation, and comprehensive explanation of the methodological procedures used within the proposed method. Our new method, called BayTTA, aims to optimize TTA using BMA and uncertainty estimation. The first component of our method involves formulating a mathematical model that applies BMA to output predictions generated from multiple transformed versions of the input data (e.g. input images). The second component of the method estimates the uncertainty of the predictions obtained from an augmented set of test images associated with image transformations or noise. Together, these components offer a comprehensive approach to enhance the robustness and reliability of ML or DL models during the testing phase. An overview of the proposed BayTTA method is presented in Fig. 3.
@@
Figure 3: An overview of the proposed BayTTA method. During the testing phase: (a) TTA generates predictions from a set of fixed augmented images, and (b) BMA is then applied to combine and aggregate these predictions by treating each unique combination as a distinct candidate model.
@@
Traditional TTA involves applying a number of data augmentation techniques, including rotation, cropping, flipping, and brightness adjustments, to the test data before making predictions. This widely used approach leverages ensemble methods, such as averaging or taking the majority vote of predictions made on augmented samples, to reduce the impact of random noise and variation in the test data. TTA usually produces stable and accurate predictions by averaging multiple augmented versions of the test dataset [11, 53].
@@
Our study implements BMA to integrate predictions obtained from TTA (i.e. predictions obtained from the same model but using different distorted images), thus deviating from the conventional averaging approach used in TTA to yield more precise and robust outcomes. Let us focus on a single image $I_{y}$ from the test dataset, where $y\in\{0,1\}$ denotes its classification label. In a scenario generating $k$ augmented versions of the test image in TTA, the input to the BMA comes from the vector ${\mathbf{x}=\{x_{1},x_{2},...,x_{k+1}\}}$, where each element corresponds to a prediction from the TTA procedure. Here, $x_{1}$ denotes the prediction derived from the original test image $I_{y}$, and $\{x_{2},x_{3},...,x_{k+1}\}$ denotes the predictions from augmented versions of this image.
@@
Rather than treating the TTA outputs as independent predictions to be averaged, we consider them as predictor variables, each with a unique combination, defining a distinct candidate model within the BMA framework. BMA involves generating various models by combining predictor variables and selecting candidate models based on likelihood. To formalize this process, we introduce $\mathcal{P(\mathbf{x})}$, a non-empty power set of $\mathbf{x}$, representing all combinations of predictor variables. Subsequently, utilizing $I\ \in\mathcal{P(\mathbf{x})}$, we select desired predictor variables from $\mathbf{x}$ for model $M_{I}$, where $\mathbf{x}_{I}$ represents the input of the respective model $M_{I}$. In the context of Bayesian model averaging logistic regression, we further elucidate the mathematical formulations that define our algorithm in the subsequent discussion.
@@
IV Experimental evaluations and discussion